Find an approximate value of the double integral below where 𝑅 is the rectangular region having



vertices (βˆ’1, 1) and (2, 3). Take a partition of 𝑅 formed by the lines π‘₯ = 0, π‘₯ = 1, and 𝑦 = 2, and take (𝑒𝑖



, 𝑣𝑖) at the



center of the 𝑖th sub region.



∬(3𝑦 βˆ’ 2π‘₯



2)𝑑𝐴



𝑅



1
Expert's answer
2022-02-01T12:35:49-0500

Question: Find an approximate value of the double integral below where 𝑅 is the rectangular region having

vertices (βˆ’1, 1) and (2, 3). Take a partition of 𝑅 formed by the lines π‘₯ = 0, π‘₯ = 1, and 𝑦 = 2, and take (𝑒𝑖, 𝑣𝑖) at the center of the 𝑖th sub region.

∬(3𝑦 βˆ’ 2π‘₯^2)𝑑𝐴

𝑅

Solution:

"\int\int_R(3y-2x^2)dA
\\=\int_{x=-1}^{x=2}\int_{y=1}^{y=3}(3y-2x^2)dxdy
\\=\int_{x=-1}^{x=2}(\dfrac{3y^2}2-2x^2y)|_{y=1}^{y=3}dx
\\=\int_{x=-1}^{x=2}[(\dfrac{3(3)^2}2-2x^2(3))-(\dfrac{3(1)^2}2-2x^2(1))]dx
\\=\int_{x=-1}^{x=2}[(\dfrac{27}2-6x^2)-(\dfrac{3}2-2x^2)]dx
\\=\int_{x=-1}^{x=2}(12-4x^2)dx
\\=(12x-\dfrac{4x^3}3)|_{x=-1}^{x=2}
\\=(12(2)-\dfrac{4(2)^3}3)-(12(-1)-\dfrac{4(-1)^3}3)
\\=(24-\dfrac{32}3)-(-12+\dfrac{4}3)
\\=36-12
\\=24\ sq.\ units"